# dragapart.github.io

### v1

@@ 0.A.2 Experiment Details
Our model was trained on 4×4$\times$4 × NVIDIA A40 GPUs with batch size $4\times 32$ and learning rate $10^{-5}$ for $150,000$ iterations using AdamW optimizer. For the last $50,000$ iterations, we sample with probability $20\%$ random texture renderings and $80\%$ regular renderings in a batch. For each batch item, we also zero out the drags $\mathcal{D}$ $10\%$ of the time to enable classifier-free guidance [23].
@@ 0.A.3 Drag-a-Move Dataset Details
While in this formulation each animation only contains parts opening, during training we randomly sample two frames, one as the reference image and the other as the ground truth image to be noised, irrespective of the order.
@@ 0.A.4 Details on Drag-Driven Moving Part Segmentation
Following existing works on semantic segmentation using image generative models, the first step is to extract the diffusion model’s internal features. Recall that at inference time the model takes an image $y$ (to be segmented) and drags $\mathcal{D}$ as input and generates another image $x\sim\mathbb{P}(x|y,\mathcal{D})$ from pure Gaussian noise. Formally, given an input image-drags pair $(y,\mathcal{D})$, we first sample a noisy latent $z_{t}$ at time step $t$ as:
@@
In our implementation, the drag conditions are injected into the self-attention layer of each UNet block. Empirically, we find using the output of the self-attention layer, instead of the whole UNet block, provides better segmentation results. Following previous works [2, 74], we resize the features at each block $l$ to the image resolution and concatenate them to create a feature pyramid.
@@ 0.B.1 Additional Qualitative Results
Refer to captionInputGenerationInputGenerationInputGeneration
Figure A: Additional Qualitative Results.
Additional results on a variety of categories can be found in Fig. A.
@@ 0.B.2 Failure Cases
Refer to captionInputGenerationInputGenerationInputGeneration
Figure B: Failure Cases. While our model is capable of generating images of complex part-level dynamics, it occasionally produces physically implausible samples.
@@
We show typical failure cases of our model in Fig. B. Notably, the model’s limitations, potentially attributed to the limited scale of the training data, manifest in several ways: i) it may incorrectly recognize different motion parts of an object (left); ii) unrealistically distort an object from an unseen category (middle); or iii) occasionally create images that lack physical plausibility (right).
@@ 0.B.3 Quantitative Comparisons of Moving Part Segmentation
Refer to captionInput ImageOursCLIPDINOSD
Figure C: Visualization of Image Features. Notably, the features extracted through our DragAPart are smooth and exhibit part-level information.
We assess our proposed moving part segmentation approach on the test split of Drag-a-Move, where part-level segmentation is annotated so the ground-truth masks are available. In Tab. A, we provide the average mask Intersection over Union (mIoU) obtained using our method and baselines of clustering alternative CLIP [50], DINO [9, 44] and SD [^5] To ensure fairness, the metric for SD is obtained by selecting the maximum from the $5$ mIoU values with diffusion time step set to 00, $200$, $500$, $800$ and $999$ respectively. [53] features. We also visualize the corresponding features (transformed through principal component analysis on foreground pixels to $3$ channels as RGB values) in Fig. C. The results underscore that the features extracted through our DragAPart enhanced part-level understanding.

### v2

@@ 0.A.2 Experiment Details
Our model was trained on 4×4$\times$4 × NVIDIA A40 GPUs with batch size $4\times 32$ and learning rate $10^{-5}$ for $150,000$ iterations using AdamW optimizer. For the last $50,000$ iterations, we sample with probability $20\%$ random texture renderings and $80\%$ regular renderings in a batch. For each batch item, we also zero out the drags $\mathcal{D}$ $10\%$ of the time to enable classifier-free guidance.
@@ 0.A.3 Drag-a-Move Dataset Details
While in this formulation, each animation only contains parts opening, during training we randomly sample two frames, one as the reference image and the other as the ground truth image to be noised, irrespective of the order.
@@ 0.A.4 Details on Drag-Driven Moving Part Segmentation
Following existing works on semantic segmentation using image generative models, the first step is to extract the diffusion model’s internal features. Recall that at inference time, the model takes an image $y$ (to be segmented) and drags $\mathcal{D}$ as input and generates another image $x\sim\mathbb{P}(x|y,\mathcal{D})$ from pure Gaussian noise. Formally, given an input image-drags pair $(y,\mathcal{D})$, we first sample a noisy latent $z_{t}$ at time step $t$ as:
@@
In our implementation, the drag conditions are injected into the self-attention layer of each UNet block. Empirically, we find using the output of the self-attention layer instead of the whole UNet block, provides better segmentation results. Following previous works [2, 71], we resize the features at each block $l$ to the image resolution and concatenate them to create a feature pyramid.
@@ 0.B.1 Additional Qualitative Results
Additional results on a variety of categories can be found in Fig. C. In instances where the conditional image features a complex, real-world background, the model adeptly “fills in” the occluded background details. This is notable given that its training exclusively utilizes images with white backgrounds. This capability is largely thanks to the strong priors of the pre-trained SD model.
@@ 0.B.2 Failure Cases
InputGenerationInputGenerationInputGeneration
Figure A: Failure Cases. While our model is capable of generating images of complex part-level dynamics, it occasionally produces physically implausible samples.
@@
We show typical failure cases of our model in Fig. A. Notably, the model’s limitations, potentially attributed to the limited scale of the training data, manifest in several ways: it may i) incorrectly recognize different motion parts of an object (left); ii) unrealistically distort an object from an unseen category (middle); or iii) occasionally create images that lack physical plausibility (right).
@@ 0.B.3 Quantitative Comparisons of Moving Part Segmentation
Input ImageOursCLIPDINOSD
Figure B: Visualization of Image Features. Notably, the features extracted through our DragAPart are smooth and exhibit part-level information.
We assess our proposed moving part segmentation approach on the test split of Drag-a-Move, where part-level segmentation is annotated so the ground-truth masks are available. In Tab. A, we provide the average mask Intersection over Union (mIoU) obtained using our method and baselines of clustering alternative CLIP [48], DINO [9, 42] and SD [^6] To ensure fairness, the metric for SD is obtained by selecting the maximum from the $5$ mIoU values with diffusion time step set to 00, $200$, $500$, $800$ and $999$ respectively. [51] features. We also visualize the corresponding features (transformed through principal component analysis on foreground pixels to $3$ channels as RGB values) in Fig. B. The results underscore that the features extracted through our DragAPart demonstrate better part-level information.
InputGenerationInputGenerationInputGeneration
Figure C: Additional Qualitative Results.